# THE MORE, THE MERRIER: CONTRASTIVE FUSION FOR HIGHER-ORDER MULTIMODAL ALIGNMENT

### v1

@@ Abstract
Learning joint representations across multiple modalities remains a central challenge in multimodal machine learning. Prevailing approaches predominantly operate in pairwise settings, aligning two modalities at a time. While some recent methods aim to capture higher-order interactions among multiple modalities, they often overlook or insufficiently preserve pairwise relationships, limiting their effectiveness on single-modality tasks. In this work, we introduce Contrastive Fusion (ConFu), a framework that jointly embeds both individual modalities and their fused combinations into a unified representation space, where modalities and their fused counterparts are aligned. ConFu extends traditional pairwise contrastive objectives with an additional fused-modality contrastive term, encouraging the joint embedding of modality pairs with a third modality. This formulation enables ConFu to capture higher-order dependencies, such as XOR-like relationships, that cannot be recovered through pairwise alignment alone, while still maintaining strong pairwise correspondence. We evaluate ConFu on synthetic and real-world multimodal benchmarks, assessing its ability to exploit cross-modal complementarity, capture higher-order dependencies, and scale with increasing multimodal complexity. Across these settings, ConFu demonstrates competitive performance on retrieval and classification tasks, while supporting unified one-to-one and two-to-one retrieval within a single contrastive framework.
@@ 1 Introduction
Learning joint representations across multiple modalities is a fundamental goal in multimodal machine learning [yuan2025survey]. Contrastive frameworks such as CLIP [radford2021learning] and ALIGN [jia2021scaling] have demonstrated that aligning paired modalities, such as images and text, enables powerful zero-shot transfer and retrieval. However, these approaches remain intrinsically pairwise, capturing only correlations between two modalities while overlooking the higher-order dependencies that emerge when three or more modalities interact [guzhov2022audioclip, jia2021scaling, girdhar2023imagebind, zhu2023languagebind, wang2024omnibind, wang2024freebind, zhou2023uni3d, xue2023ulip, xue2024ulip, wang2023connecting, zhang2024extending]. Beyond this structural limitation, multimodal models often exhibit modality competition [huang2022modality], where one modality dominates prediction and suppresses others, particularly when signal strength or reliability differs across modalities. For instance, strong visual cues may overshadow weaker auditory ones, leading the joint representation to rely disproportionately on a single source. Many real-world settings, however, require reasoning over multiple complementary cues. A song, for example, emerges from the interplay between melody and lyrics, a 3D design may be defined jointly by a sketch and a text description, and a scene can be better understood when visual and auditory signals are considered together. These challenges motivate a broader question:
Can a contrastive learning framework capture not only pairwise alignments but also higher-order, synergistic dependencies among modalities?
To address this, we propose Contrastive Fusion (ConFu), a framework that unifies pairwise and higher-order multimodal supervision within a single contrastive objective. ConFu constructs fused representations of all subsets of modalities and aligns all combinations of unimodal and fused representations (Fig. 2). Unlike previous methods that primarily target specific one-to-one (11) [guzhov2022audioclip, girdhar2023imagebind], two-to-one (21) [saporta2024contrasting, cicchetti2025triangle], or both within predominantly text-centered objectives [cicchetti2024gramian, chen2023vast], ConFu accounts for all pairwise and higher-order modality combinations within a unified objective (Fig. 1).
Evaluating such models is further complicated by the scarcity of publicly available datasets that naturally include three or more co-occurring modalities. To address this, we construct a bird-centric dataset where image, audio, and text provide complementary but related information. This setting enables a systematic examination of higher-order alignment under zero-shot and few-shot regimes.
Overall, our main contributions are as follows:
@@ 2 Related Work
Most contrastive frameworks in multimodal learning operate on paired modalities. Recent extensions, categorized as jointly-aligned or pivot-aligned, expand this formulation to multiple modalities, yet they model higher-order relations only implicitly through shared objectives.
Jointly-aligned models.
@@
Methods such as AudioCLIP [guzhov2022audioclip], VALOR [chen2023valor], and VAST [chen2023vast] extend CLIP-style contrastive learning to three or more modalities by training on datasets where all signals co-occur. AudioCLIP employs independent contrastive objectives for each modality pair, while VALOR introduces a multimodal grouping alignment loss coupled with caption generation to connect multiple modalities through text. VAST combines contrastive and generative losses to unify video, audio, and subtitles in a shared representation. Other architectures such as mPLUG-2 [xu2023mplug], UMT [liu2022umt], and the Everything-at-Once Fusion Transformer [lim2021temporal] employ modular decoders or fusion transformers for multimodal understanding and generation. While these methods learn representations across multiple modalities, their objectives are typically limited to pairwise alignments or remain primarily text-centric, with broader multimodal relationships emerging implicitly from shared training signals rather than through explicit higher-order alignment.
Pivot-aligned models.
@@
An alternative strategy uses a pivot modality as a common reference space. Each non-pivot modality is aligned to this anchor, enabling scalable training without requiring full multimodal co-occurrence. For example, ImageBind [girdhar2023imagebind] aligns diverse modalities (audio, depth, IMU) to an image-based embedding space, while LanguageBind [zhu2023languagebind] uses text as the pivot. Later variants such as OmniBind [wang2024omnibind], FreeBind [wang2024freebind], and Ex-MCR [zhang2024extending] extend this paradigm through pseudo-pair generation, offering an alternate data-efficient mechanism. Similarly, Uni3D [zhou2023uni3d], ULIP [xue2023ulip], and ULIP-2 [xue2024ulip] project new modalities into a frozen CLIP embedding space,mapping new modalities to CLIP’s pretrained space without updating its parameters. While this pivot-based design improves scalability, it remains inherently pairwise: all interactions are mediated through the pivot, preventing direct modeling of dependencies among non-pivot modalities. As a result, the learned embedding space captures independent pairwise relations rather than true higher-order structure.
Modeling Higher-Order Dependencies.
@@
Recent work in self-supervised representation learning has attempted to model relationships between more than two modalities through specialized loss formulations. SYMILE [saporta2024contrasting] leverages total correlation to integrate complementary cues, TRIANGLE [cicchetti2025triangle] uses the area spanned by three embedding vectors as a measure of similarity, and GRAM [cicchetti2024gramian] measures cross-modal similarity for modalities through the Gramian volume spanned by their vectors. While these methods jointly model multiple modalities, SYMILE and TRIANGLE require all modalities to be present at inference time, making them incompatible with standard 11 retrieval.
@@ 3.1 Proposed Multimodal Contrastive Framework
To model higher-order dependencies , we introduce a fusion network that combines encoded features:
@@
In practice, both density ratio estimators and are implemented as temperature-scaled dot-product similarities. The higher-order objective aggregates all such triplets:
@@ 3.2 Theoretical Motivation
The total correlation (TC) [watanabe1960information], also known as multi-information [studeny1998multiinformation], quantifies the joint statistical dependency among multiple random variables as the Kullback–Leibler divergence between their joint distribution and the product of marginals, . Equivalently,
@@
Each MI term in Eq. 10 is approximated using the InfoNCE lower bounds [oord2018representation]. Specifically, the pairwise and higher-order terms satisfy
@@
so that maximizing total correlation reduces to minimizing the InfoNCE losses across all pairwise and higher-order relations. When each density ratio function or is sufficiently expressive and approximates the optimal InfoNCE critic [oord2018representation] (i.e., the log–density ratio between the joint and the product of marginals) the bound becomes tight as , recovering the true .
@@ 3.3 Compare with other higher-order methods
Recent approaches such as SYMILE, GRAM, and TRIANGLE adopt a single joint alignment objective across all modalities, without separating pairwise from higher-order dependencies. SYMILE maximizes a total-correlation bound using a multilinear similarity , GRAM enforces global alignment via the parallelotope volume induced by the Gram matrix, and TRIANGLE specializes this idea by introducing a triangle-area–based measure of three-way similarity. In contrast, our method decomposes total correlation into interpretable information-theoretic factors, averaging over modality permutations the sum of pairwise and higher-order dependencies, thus preserving holistic expressivity while enabling structured optimization and reliable performance with any subset of modalities.
@@ 4 Bird-MML: A Multimodal Dataset for Audio-Visual Complementarity
Despite substantial progress in multimodal learning, there remains a lack of standardized and publicly available realistic datasets that capture cross-modal relationships that are suitable for evaluating both pairwise and higher-order multimodal alignment.
@@
Bird species provide an ideal testbed for such multimodal evaluation, as their identity and behavior are naturally expressed through both visual and acoustic cues. In many cases, fine-grained recognition further depends on complementary information such as visual appearance for morphology and audio for vocalization, making the domain well suited for studying multimodal interaction.
Existing Bird Datasets.
@@
SSW60 [van2022exploring] offers unaligned images and audio as well as video for 60 species, VB100 [ge2016exploiting] focuses on fine-grained video recognition for 100 species. Moreover, iNatSounds [chasmai2024inaturalist] contains large-scale bioacoustic recordings without visual counterparts, and lastly, BioTrove-Train [yang2024biotrove] includes millions of images across diverse species with uneven per-class distributions. Although those datasets provide valuable resources, joint combination of images, audio, and text in a form suitable for multimodal pretraining and evaluation is missing.
Dataset Overview.
@@
To bridge this gap, we construct Bird-MML, a synthetic multimodal bird dataset integrating artificially paired images, audio, and text for 150 species, including all those present in SSW60 and VB100. The dataset is designed for scalable pretraining and cross-modal learning, with balanced representation across species. A detailed overview of the data generation pipeline is provided in Appendix D.1. Images are sourced from the iNaturalist open dataset, restricted to research-grade, Creative Commons–licensed observations. For each species, 1,000 samples are randomly selected.
Audio and Text Generation.
@@
Audio recordings are collected from Xeno-Canto [xenocanto], segmented into 10-second clips, and zero-padded when necessary. In cases of limited availability, existing clips are reused to maintain uniform sample counts. Metadata such as call type, sex, and life stage are preserved. Text descriptions are generated by combining three complementary sources: image captions from InstructBLIP2 [dai2023instructblip], audio metadata, and short summaries from corresponding Wikipedia entries. These elements are fused using google/gemma-2-2b-it [gemma_2024] to yield a single multimodal caption per instance.
Dataset Characteristics.
@@
While generated captions may contain minor factual noise, they effectively capture species-relevant attributes and contextual cues useful for representation learning. The resulting dataset contains 149,681 triplets (image, audio, text), each grounded by species identity and visual–acoustic correspondence. Approximately 43% of species required limited audio reuse due to data scarcity. Examples of artificially aligned triplets are shown in Appendix D.1.The dataset maintains balanced per-species coverage and is well-suited for multimodal pretraining. Downstream evaluations are performed on naturally paired video-audio data to assess real-world cross-modal performance.
Ethical Considerations.
@@
All data were collected from publicly available sources under Creative Commons licenses. Ownership and attribution remain with original contributors. Automatically generated captions may contain minor factual errors or biases and should not be interpreted as verified biological information. The dataset is released solely for non-commercial research use.
5 Experimental Results
Our evaluation covers multimodal benchmarks across biodiversity (SSW60 [van2022exploring], VB100 [ge2016exploiting]), affective (MOSI [zadeh2016mosi], UR-FUNNY [hasan2019ur], and MUStARD [castro2019towards] included in MultiBench [liang2021multibench]), and multimedia (AV-MNIST [vielzeuf2018centralnet]) domains, encompassing both real-world and synthetic datasets.
These datasets present different levels of cross-modal difficulty, enabling holistic evaluation of alignment, representation quality, and scalability. We compare against representative contrastive and higher-order multimodal approaches: (i) a Bi-modal CLIP (referenced as CLIP from now on) trained with standard two-modality contrastive loss, (ii) a Tri-modal CLIP (referenced as Tri-CLIP from now on) trained with pairwise objectives across all modalities, (iii) SYMILE [saporta2024contrasting] which models higher-order dependencies via total correlation, (iv) TRIANGLE [cicchetti2025triangle] and GRAM [cicchetti2024gramian], which use geometric similarity measures among 2 or more modalities. Experiments cover (i) zero-shot classification, (ii) 11 and 21 retrieval, and (iii) linear-probe classification under few-shot and full-data regimes. Details about the experimental setup and evaluation can be found in Appendix A.
5.1 AV-MNIST
We evaluate our framework on the AV-MNIST dataset, a controlled audio–visual benchmark. The dataset contains paired audio spectrograms augmented with randomly sampled background noise from ESC-50 [piczak2015dataset] and images of MNIST digits with degraded visual features. We additionally introduce a text modality derived from class labels using natural language templates such as ‘the digit is a {class}”, ‘a photo of a {class}”, and ‘a picture of a {class}”. As shown in Table 1, our approach achieves the highest performance across all baselines, with Symile trailing closely behind. Fusing audio and visual inputs yields an 8% improvement over the strongest unimodal baseline, demonstrating the benefit of complementary multimodal supervision. Moreover, even when evaluated on a single modality (vision), our model surpasses the corresponding unimodal baseline by 1.5%, indicating that multimodal training enhances individual modality representations.
@@ 5.1 AV-MNIST
| Symile [saporta2024contrasting] | - | - | 70.9 0.3 |
| GRAM [cicchetti2024gramian] | 9.8 0.4 | 63.9 0.8 | 64.4 0.5 |
| TRIANGLE [cicchetti2025triangle] | - | - | 64.9 0.3 |

### v2

@@ Abstract
Learning joint representations across multiple modalities remains a central challenge in multimodal machine learning. Prevailing approaches predominantly operate in pairwise settings, aligning two modalities at a time. While some recent methods aim to capture higher-order interactions among multiple modalities, they often overlook or insufficiently preserve pairwise relationships, limiting their effectiveness on single-modality tasks. In this work, we introduce Contrastive Fusion (ConFu), a framework that jointly embeds both individual modalities and their fused combinations into a unified representation space, where modalities and their fused counterparts are aligned. ConFu extends traditional pairwise contrastive objectives with an additional fused-modality contrastive term, encouraging the joint embedding of modality pairs with a third modality. This formulation enables ConFu to capture higher-order dependencies, such as XOR-like relationships, that cannot be recovered through pairwise alignment alone, while still maintaining strong pairwise correspondence. We evaluate ConFu on synthetic and real-world multimodal benchmarks, assessing its ability to exploit cross-modal complementarity, capture higher-order dependencies, and scale with increasing multimodal complexity. Across these settings, ConFu demonstrates competitive performance on retrieval and classification tasks, while supporting unified one-to-one and two-to-one retrieval within a single contrastive framework. We release our code and dataset at https://github.com/estafons/confu.
@@ 1 Introduction
Learning joint representations across multiple modalities is a fundamental goal in multimodal machine learning [44]. Contrastive frameworks such as CLIP [26] and ALIGN [16] have demonstrated that aligning paired modalities, such as images and text, enables powerful zero-shot transfer and retrieval. However, these approaches remain intrinsically pairwise, capturing only correlations between two modalities while overlooking the higher-order dependencies that emerge when three or more modalities interact [11, 16, 10, 48, 35, 34, 47, 41, 42, 36, 46]. Beyond this limitation, multimodal models often exhibit modality competition [14], where one modality dominates and suppresses others when signal strength or reliability differs. For instance, strong visual cues may overshadow weaker auditory ones, biasing the joint representation toward a single source. Yet many real-world settings require reasoning over complementary cues: a song emerges from melody and lyrics, a 3D design from sketch and text, and a scene from both visual and auditory signals. These challenges motivate a broader question:
To address this, we propose Contrastive Fusion (ConFu), a framework that unifies pairwise and higher-order multimodal supervision within a single contrastive objective. ConFu constructs fused representations of all subsets of modalities and aligns all combinations of unimodal and fused representations (Fig. 2). Unlike previous methods that primarily target specific one-to-one (11) [11, 10], two-to-one (21) [27, 5], or both within predominantly text-centered objectives [6, 4], ConFu accounts for all pairwise and higher-order modality combinations within a unified objective (Fig. 1).
Evaluating such models is further complicated by the scarcity of publicly available datasets that naturally include three or more co-occurring modalities. To address this, we construct a bird-centric dataset where image, audio, and text provide complementary but related information. This setting enables a systematic examination of higher-order alignment under zero-shot and few-shot regimes. Overall, our main contributions are as follows:
@@ 2 Related Work
Jointly-aligned models. Methods such as AudioCLIP [11], VALOR [3], and VAST [4] extend CLIP-style contrastive learning to three or more modalities by training on datasets where all signals co-occur. AudioCLIP employs independent contrastive objectives for each modality pair, while VALOR introduces a multimodal grouping alignment loss coupled with caption generation to connect multiple modalities through text. VAST combines contrastive and generative losses to unify video, audio, and subtitles in a shared representation. Other architectures such as mPLUG-2 [40], UMT [22], and the Everything-at-Once Fusion Transformer [21] employ modular decoders or fusion transformers for multimodal understanding and generation. While these methods learn representations across multiple modalities, their objectives are typically limited to pairwise alignments or remain primarily text-centric, with broader multimodal relationships emerging implicitly from shared training signals rather than through explicit higher-order alignment.
@@
Pivot-aligned models. An alternative strategy uses a pivot modality as a common reference space. Each non-pivot modality is aligned to this anchor, enabling scalable training without requiring full multimodal co-occurrence. For example, ImageBind [10] aligns diverse modalities (audio, depth, IMU) to an image-based embedding space, while LanguageBind [48] uses text as the pivot. Later variants such as OmniBind [35], FreeBind [34], and Ex-MCR [46] extend this paradigm through pseudo-pair generation, offering an alternate data-efficient mechanism. Similarly, Uni3D [47], ULIP [41], and ULIP-2 [42] project new modalities into a frozen CLIP embedding space,mapping new modalities to CLIP’s pretrained space without updating its parameters. While this pivot-based design improves scalability, it remains inherently pairwise: all interactions are mediated through the pivot, preventing direct modeling of dependencies among non-pivot modalities. As a result, the learned embedding space captures independent pairwise relations rather than true higher-order structure.
@@
Modeling Higher-Order Dependencies. Recent work in self-supervised representation learning has attempted to model relationships between more than two modalities through specialized loss formulations. Symile [27] leverages total correlation to integrate complementary cues, TRIANGLE [5] uses the area spanned by three embedding vectors as a measure of similarity, and GRAM [6] measures cross-modal similarity for modalities through the Gramian volume spanned by their vectors. While these methods jointly model multiple modalities, Symile and TRIANGLE require all modalities to be present at inference time, making them incompatible with standard 11 retrieval.
@@ 3.1 Proposed Multimodal Contrastive Framework
To model higher-order dependencies , we introduce a fusion network with encoded features:
@@
In practice, both density ratio estimators in Eq. 2 and in Eq. 5 are implemented as temp.-scaled dot-product similarities. The higher-order objective aggregates all such triplets:
@@ 3.2 Theoretical Motivation
The total correlation (TC) [37], also known as multi-information [28], quantifies the joint statistical dependency among multiple random variables as the Kullback–Leibler divergence between their joint distribution and the product of marginals, . Equivalently,
@@
Each MI term in Eq. 10 is approximated using the InfoNCE lower bounds [23]. Specifically, the pairwise and higher-order terms satisfy
@@
so that maximizing total correlation reduces to minimizing the InfoNCE losses across all pairwise and higher-order relations. When each density ratio function or is sufficiently expressive and approximates the optimal InfoNCE critic [23] (i.e., the log–density ratio between the joint and the product of marginals) the bound becomes tight as , recovering the true .
@@ 3.3 Compare with other higher-order methods
Recent approaches such as Symile, GRAM, and TRIANGLE adopt a single joint alignment objective across all modalities, without separating pairwise from higher-order dependencies. Symile maximizes a total-correlation bound using a multilinear similarity , GRAM enforces global alignment via the parallelotope volume induced by the Gram matrix, and TRIANGLE specializes this idea by introducing a triangle-area–based measure of three-way similarity. In contrast, our method decomposes total correlation into interpretable information-theoretic factors, averaging over modality permutations the sum of pairwise and higher-order dependencies, thus preserving holistic expressivity while enabling structured optimization and reliable performance with any subset of modalities. Unlike prior approaches that use a single critic to implicitly model dependencies, we factorize them at the loss level, enabling independent control of pairwise and higher-order supervision while preserving InfoNCE stability and achieving balanced performance within a single model.
@@ 4 Bird-MML: A Multimodal Dataset for Audio-Visual Complementarity
Despite substantial progress in multimodal learning, there remains a lack of standardized and publicly available realistic datasets that capture cross-modal relationships that are suitable for evaluating both pairwise and higher-order multimodal alignment. Bird species provide an ideal testbed for such multimodal evaluation, as their identity and behavior are naturally expressed through both visual and acoustic cues. In many cases, fine-grained recognition further depends on complementary information such as visual appearance for morphology and audio for vocalization, making the domain well suited for studying multimodal interaction.
@@
Existing Bird Datasets. SSW60 [30] offers unaligned images and audio as well as video for 60 species, VB100 [9] focuses on fine-grained video recognition for 100 species. Moreover, iNatSounds [2] contains large-scale bioacoustic recordings without visual counterparts, and lastly, BioTrove-Train [43] includes millions of images across diverse species with uneven per-class distributions. Although those datasets provide valuable resources, a joint combination of images, audio, and text in a form suitable for multimodal pretraining and evaluation is missing.
@@
Dataset Overview. To bridge this gap, we construct Bird-MML, a synthetic multimodal bird dataset integrating artificially paired images, audio, and text for 150 species, including all those present in SSW60 and VB100. The dataset is designed for scalable pretraining and cross-modal learning, with balanced representation across species. A detailed overview of the data generation pipeline is provided in Appendix D.1. Images are sourced from the iNaturalist open dataset, restricted to research-grade, Creative Commons–licensed observations. For each species, 1,000 samples are randomly selected.
@@
Audio and Text Generation. Audio recordings are collected from Xeno-Canto [39], segmented into 10-second clips, and zero-padded when necessary. In cases of limited availability, existing clips are reused to maintain uniform sample counts. Metadata such as call type, sex, and life stage are preserved. Text descriptions are generated by combining three complementary sources: image captions from InstructBLIP2 [7], audio metadata, and short summaries from corresponding Wikipedia entries. These elements are fused using google/gemma-2-2b-it [29] to yield a single multimodal caption per instance.
@@
Dataset Characteristics. While generated captions may contain minor factual noise, they effectively capture species-relevant attributes and contextual cues useful for representation learning. The resulting dataset contains 149,681 triplets (image, audio, text), each grounded by species identity and visual–acoustic correspondence. Approximately 43% of species required limited audio reuse due to data scarcity. Examples of artificially aligned triplets are shown in Appendix D.1.The dataset maintains balanced per-species coverage and is well-suited for multimodal pretraining. Downstream evaluations are performed on naturally paired video-audio data to assess real-world cross-modal performance.
@@
Ethical Considerations. All data were collected from publicly available sources under Creative Commons licenses. Ownership and attribution remain with original contributors. Automatically generated captions may contain minor factual errors or biases and should not be interpreted as verified biological information. The dataset is released solely for non-commercial research use.
@@
| Symile [27] | - | - | 70.9 0.3 |
| GRAM [6] | 9.8 0.4 | 63.9 0.8 | 64.4 0.5 |
| TRIANGLE [5] | - | - | 64.9 0.3 |
@@
5 Experimental Results
Our evaluation covers multimodal benchmarks across biodiversity (SSW60 [30], VB100 [9]), affective (MOSI [45], UR-FUNNY [12], and MUStARD [1] included in MultiBench [20]), and multimedia (AV-MNIST [31]) domains, encompassing both real-world and synthetic datasets.
These datasets present different levels of cross-modal difficulty, enabling holistic evaluation of alignment, representation quality, and scalability. We compare against representative contrastive and higher-order multimodal approaches: (i) a Bi-modal CLIP (referenced as CLIP from now on) trained with standard two-modality contrastive loss, (ii) a Tri-modal CLIP (referenced as Tri-CLIP from now on) trained with pairwise objectives across all modalities, (iii) Symile [27] which models higher-order dependencies via total correlation, (iv) TRIANGLE [5] and GRAM [6], which use geometric similarity measures among 2 or more modalities. Experiments cover (i) zero-shot classification, (ii) 11 and 21 retrieval, and (iii) linear-probe classification under few-shot and full-data regimes. Details about the experimental setup and evaluation can be found in Appendix A.
5.1 AV-MNIST
We evaluate our framework on the AV-MNIST dataset, which contains paired audio spectrograms augmented with randomly sampled background noise from ESC-50 [25] and images of MNIST digits with degraded visual features. We additionally introduce a text modality derived from class labels using natural language templates such as ‘the digit is a {class}”, ‘a photo of a {class}”, and ‘a picture of a {class}”. As provided in Table 1, our approach achieves the highest performance across all baselines, with Symile trailing closely behind. Fusing audio and visual inputs yields an 8% improvement over the strongest unimodal baseline, demonstrating the benefit of complementary multimodal supervision. Moreover, even when evaluated on a single modality (vision), our model surpasses the corresponding unimodal baseline by 1.5%, indicating that multimodal training enhances individual modality representations.